OpenStudy (anonymous):

Help with algebra 2? Best answer gets a medal.

OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -1 /2 and b = 1

OpenStudy (anonymous):

we have : y=mx+d So have :y=-1/2+1 Just this ;) Got it ?:)

OpenStudy (anonymous):

Yup! Thanks!

OpenStudy (anonymous):

Welcome ;)